Solve the following triangles. Identify the case # in each given
triangle.

3.) Given parts:

a= 4 b= 8 A=30°
1
Expert's answer
2020-11-09T20:08:57-0500

A=30"\degree" , a=4 , b=8


step 1. use the law of cosines to find the side

c2 = b2 + a2 -2ba cosA

c2 = 82 + 42 - 2(8)(4) cos30"\degree"

c2= 64 + 16 + 64 (0.86)

c2 = 80 + 55.04

c2 = 135

c = 11.61

step 2. use the law of sin to find the opposite angle

"\frac{b}{sinB}" = "\frac{a}{sinA}"


"\frac{8}{SinB}" = "\frac{4}{sin30}"


4sinB =8sin30"\degree"

sinB = 2 sin30"\degree"

sinB = 1

B = 90"\degree"

step 3. find the third angle

C = 180"\degree" - A - B

C = 180"\degree" - 30"\degree" - 90"\degree"

C = 60"\degree"

the solution is C= 60"\degree" B = 90"\degree" and c = 11.61